Operations Report – May 2nd

Crew 193 Operations Report 02May2018

SOL: 10

Name of person filing report: David Attig

Non-nominal systems: Two piece suits still need straps adjusted for neck ring. Generator/inverter/charger. One piece suit #10.

Notes on non-nominal systems: The generator still appears to not be charging the batteries at night (only powering the HAB). Shannon is having a look, this should not affect our operations (particularly with the days being so long now).

Generator (hours run): 7:00 pm to 7:55 am.

Solar— SOC 60 % (Before generator is run at night)

Diesel – 70 %

Propane – 60 psi.

ATV Fuel (Ethanol free gasoline) – 0 gallons.

Water (trailer) – 220 gallons.

Water (static) – 290 gallons

Trailer to Static Pump used – no

Water (loft) – Static to Loft Pump used – yes

Water Meter: 134505.5

Toilet tank emptied: yes
Date last emptied: May 2

ATV’s (Honda, 300, 350.1, 350.2, 350.3): No

ATV Oil Checked?: No

ATV Oil Added?: No

ATV Fuel Used: 0 Gallons

# Hours the ATVs were Used today: 0

Notes on ATVs: 350.2 is out of service, due to a previous issue. Honda has a hole in the seat.

Deimos rover used: No
Hours: 124.8
Beginning charge: 100%
Ending charge: N/A %
Currently charging: No

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: Unknown
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

Spirit rover used: No
Hours: 35.0
Beginning charge: 100 %
Ending charge: N/A %
Currently charging: No

Opportunity rover used: No
Hours: 34.2
Beginning charge: 100 %
Ending charge: N/A %
Currently charging: Yes

Curiosity rover used: No
Hours: 29.8
Beginning charge: 100 %
Ending charge: N/A %
Currently charging: Yes

Hab Car used?: No
Where was Hab Car used?: N/A
Why was Hab Car used?: N/A

General notes and comments: Muddy

Summary of internet: We’re out, same as normal.

Summary of suits and radios: One piece suit #10 is repaired. Still work to do on two piece suits. One suit charger repaired, repair of other suit chargers still in progress.

Summary of Hab operations: Muddily nominal

Summary of GreenHab operations: Muddy

Summary of ScienceDome operations: Muddy

Summary of RAM operations: None today, Muddy

Summary of health and safety issues: Muddy

Questions, concerns and requests to Mission Support: Less mud. And an escalator.

Copyright © The Mars Society. All rights reserved. | Main Site